# SAINT: Attention-Based Policies for Discrete Combinatorial Action Spaces

### v2

@@ Abstract
The combinatorial structure of many real-world action spaces leads to exponential growth in the number of possible actions, limiting the effectiveness of conventional reinforcement learning algorithms. Recent approaches for combinatorial action spaces impose factorized or sequential structures over sub-actions, failing to capture complex joint behavior. We introduce the Sub-Action Interaction Network using Transformers (SAINT), a novel policy architecture that represents multi-component actions as unordered sets and models their dependencies via self-attention conditioned on the global state. SAINT is permutation-invariant, sample-efficient, and compatible with standard policy optimization algorithms. In 20 distinct combinatorial environments across three task domains, including environments with nearly 17 million joint actions, SAINT consistently outperforms strong baselines.  [^1] Code is available at https://github.com/matthewlanders/SAINT
@@ 1 Introduction
Traditional RL methods model discrete action spaces with a flat categorical policy, but this becomes intractable in combinatorial settings where the number of actions scales as $\prod_{d=1}^{A}m_{d}$ for $A$ sub-action dimensions with $m_{d}$ choices each. To mitigate this combinatorial explosion, existing approaches (Dulac-Arnold et al., 2015; Pierrot et al., 2021; Tavakoli et al., 2018; Zhang et al., 2018) rely on simplifying assumptions that constrain the representational capacity of the policy class. One family of methods (Pierrot et al., 2021; Tavakoli et al., 2018) factorizes the policy as $\pi(\mathbf{a}\mid s)=\prod_{i}\pi_{i}(a_{i}\mid s)$, which cannot represent interactions between sub-actions. Another class of approaches (Zhang et al., 2018) imposes a fixed autoregressive order, specifying a policy class of distributions of the form $\pi(\mathbf{a}|s)=\prod_{i}\pi_{i}(a_{i}|s,a_{<i})$. This introduces an arbitrary sequence over sub-actions, breaking permutation invariance and impairing learning when the imposed order misaligns with the true dependency structure. Many real-world tasks violate both the independence and fixed-order assumptions. In healthcare, for example, drug combinations can exhibit complex interaction effects — treatments may be safe individually but harmful together, motivating permutation-invariant models for combinatorial action spaces. Our work targets precisely these settings: combinatorial action spaces wherein sub-action indexing is arbitrary or only weakly meaningful, and the fundamental structure lies in sub-action interactions rather than in any prescribed ordering.
@@
We evaluate SAINT on challenging benchmark tasks, which exhibit both state-independent and state-dependent sub-action dependencies, including traffic light control (Zhang et al., 2019), navigation (Landers et al., 2024), and discretized MuJoCo locomotion tasks (Towers et al., 2024). Our results demonstrate that by learning a more expressive representation of the action space’s internal structure, SAINT consistently outperforms strong factorized and autoregressive baselines, scaling to environments with nearly 17 million discrete actions. Targeted ablations validate the role of state conditioning and show that the additional cost of modeling dependencies is often offset by substantial gains in sample efficiency. Together, these findings establish that learning explicit representations of sub-action interactions is a practical and scalable approach to decision-making in complex combinatorial domains.
@@ 2 Related Work
Combinatorial action spaces arise naturally in sequential decision problems such as traffic signal control, games, and resource allocation. Prior work has introduced task-specific architectures, imposed domain-specific assumptions, or exploited problem-specific structure (Bello et al., 2016; Chen et al., 2023; Delarue et al., 2020; He et al., 2015; 2016; Nazari et al., 2018; Zahavy et al., 2018; Farquhar et al., 2020). Such methods typically lack generality and require manual design effort. A parallel body of work addresses continuous control problems by discretizing the action space (Barth-Maron et al., 2018; Metz et al., 2017; Tang and Agrawal, 2020; Van de Wiele et al., 2020), which contrasts with our focus on inherently discrete action spaces with combinatorial structure.
@@
A number of general-purpose architectures have been developed to scale RL to large combinatorial action spaces. One strategy reduces complexity by assuming conditional independence across sub-actions (Pierrot et al., 2021; Tavakoli et al., 2018), while another imposes an autoregressive order (Zhang et al., 2018). These approaches improve tractability but either ignore dependencies among sub-actions or introduce arbitrary orderings that break permutation symmetry. Retrieval-based methods such as Wolpertinger (Wol-DDPG) (Dulac-Arnold et al., 2015) scale to large discrete spaces by embedding and pruning candidate actions, but similarly fail to capture the joint structure of unordered sub-actions (Chen et al., 2023). These limitations motivate architectures that can represent dependencies across sub-actions while preserving permutation invariance.
Transformers for Action Representation
@@
Efforts to use Transformers for action space modeling have largely focused on sequential representations. RT-1 (Brohan et al., 2022) and RT-2 (Zitkovich et al., 2023) tokenize robot control trajectories and decode action tokens autoregressively, while Q-Transformer (Chebotar et al., 2023) autoregresses across action dimensions within a timestep. Learned tokenizers such as FAST (Pertsch et al., 2025) compress high-frequency control signals into vocabularies for vision–language–action (VLA) training. Extensions of trajectory-based models (Chen et al., 2021; Shang et al., 2022), interleave state and action tokens or design state-aware tokenizations to handle multi-component actions. Together, these methods establish the efficacy of Transformers for action encoding, but have been mostly designed for continuous control tasks in offline RL or VLA settings that are not directly comparable to the combinatorial domains we study. Moreover, they largely rely on sequential decompositions that impose arbitrary order and obscure permutation symmetry.
@@ 3 Preliminaries
Attention is a general computational primitive that allows a model to selectively aggregate information from a set of inputs based on learned relevance scores (Vaswani et al., 2017). Formally, given a set of queries $Q\in\mathbb{R}^{n_{q}\times d}$, keys $K\in\mathbb{R}^{n_{k}\times d}$, and values $V\in\mathbb{R}^{n_{k}\times d}$, the scaled dot-product attention computes output $\mathrm{Attn}(Q,K,V)=\text{softmax}(QK^{\top}/\sqrt{d})V$. Multi-head self-attention extends this by learning multiple independent projections and aggregating their outputs, enabling the model to capture different types of interactions in parallel. Transformers, which stack layers of multi-head self-attention with feedforward components, have been widely adopted in domains requiring flexible modeling of structured dependencies.
@@ 4.1 State Conditioning
Because sub-action identity alone is insufficient for decision-making, SAINT augments each base embedding $\mathbf{e}_{i}$ with information from the global state $\mathbf{s}\in\mathbb{R}^{d_{s}}$, enabling dependencies to be modeled in a state-aware manner. While several conditioning mechanisms such as cross-attention or concatenation are possible, we adopt Feature-wise Linear Modulation (FiLM) (Perez et al., 2018), which we found to be effective and parameter-efficient (see Appendix C). Notably, FiLM preserves the $d$-dimensional width of each sub-action embedding, introducing no additional projection dimensions. Its prior success in incorporating state information in RL (Brohan et al., 2022) further supports this choice.
An MLP $g:\mathbb{R}^{d_{s}}\to\mathbb{R}^{2d}$ processes the global state $\mathbf{s}$ once to produce FiLM parameters $(\bm{\gamma},\bm{\beta})=g(\mathbf{s})$, which are then applied uniformly to all sub-action embeddings via an affine transformation:
$\tilde{\mathbf{e}}_{i}=\bm{\gamma}\odot\mathbf{e}_{i}+\bm{\beta}\;.$
@@ 4.3 Action Decoding
1:Initialize SAINT policy $\pi_{\theta}$ and value network $V_{\phi}$
@@
3:  Collect a batch of transitions $(\mathbf{s}_{t},\mathbf{a}_{t},r_{t},\text{done}_{t})$ by executing $\pi_{\theta}$
4:  Compute return $R_{t}$ and weighting term $w_{\Phi}(\mathbf{s}_{t},\mathbf{a}_{t},R_{t})$ for each transition
5:  Compute policy log-probabilities $\ell_{t}\leftarrow\textsc{LogProbs}(\mathbf{s}_{t},\mathbf{a}_{t})$ $\triangleright$ See function below
6:  Update policy $\theta$ by ascending the objective $\mathbb{E}_{t}[w_{\Phi}(\mathbf{s}_{t},\mathbf{a}_{t})\cdot\ell_{t}]$
7:  Update value function $\phi$ by descending the loss $\mathbb{E}_{t}[(V_{\phi}(\mathbf{s}_{t})-R_{t})^{2}]$
@@
11:  Get sub-action embeddings $\mathbf{E}\leftarrow[\text{Embed}(1),\dots,\text{Embed}(A)]^{\top}$
@@
14:  Get logits for each sub-action $\text{Logits}_{i}\leftarrow f_{i}(\mathbf{X}_{[:,i]})$ for $i=1,\dots,A$
15:  Compute log-probabilities $\log\mathbf{P}_{[:,i]}\leftarrow\log\pi_{i}(\mathbf{A}_{\text{taken}[:,i]}\mid\text{Logits}_{i})$ for $i=1,\dots,A$
@@ 4.4 Compatibility with RL Algorithms
Incompatibility arises when the actor objective requires global operations over the entire combinatorial action space, such as $\mathbb{E}_{\mathbf{a}^{\prime}\sim\pi*\theta}[Q_{\Phi}(s,\mathbf{a}^{\prime})]$ or $\max_{\mathbf{a}^{\prime}}Q_{\Phi}(s,\mathbf{a}^{\prime})$. These operations are computationally intractable unless $Q_{\Phi}$ is factorized; however, this changes the structure of the actor target, decomposing it into uncoordinated per-dimension terms and discarding cross-dimensional structure. This breaks alignment with SAINT’s objective of modeling dependencies among sub-actions. SAINT’s learning procedure is provided in Algorithm 1.
@@ 5 Experimental Evaluation
Our experiments evaluate the efficacy of different action representations for modeling complex sub-action interactions across three regimes: (1) primarily state-independent interactions, (2) state-dependent interactions, and (3) weak interactions with complex dynamics. Results are presented in Sections 5.15.2, and 5.3, respectively. Section 5.4 evaluates SAINT in the offline RL setting. Section 5.5 analyzes key architectural choices, quantifying the trade-off between representational power and computational cost, and robustness to Transformer structural parameters.
@@
We compare SAINT to four baselines reflecting the standard representational assumptions for combinatorial action spaces: (1) a factorized policy (Tavakoli et al., 2018), assuming fully independent sub-actions; (2) an autoregressive model (Zhang et al., 2018), imposing a fixed sequential order; (3) Wol-DDPG (Dulac-Arnold et al., 2015), using a continuous embedding; and (4) a flat RL algorithm, which learns a monolithic representation of the full action space without exploiting its combinatorial structure. Results are averaged over five random seeds.
@@
These four baselines instantiate the dominant structural assumptions used to scale RL to large combinatorial action spaces. The flat policy corresponds to a monolithic model that ignores compositional structure and treats each joint action as an atomic symbol. The factorized policy enforces independent per-dimension decisions, preventing it from modeling necessary coordination between sub-actions. The autoregressive model imposes a fixed sequential ordering over sub-actions, which can be misaligned with the true, permutation-invariant dependency structure. Wol-DDPG embeds each joint action as a single continuous vector, collapsing the internal structure needed to capture interactions among sub-actions. Our experiments in Sections 5.1-5.4 are designed to test whether these structural assumptions remain sufficient when sub-action indexing is arbitrary or only weakly meaningful, or whether a set-based alternative such as SAINT is required.
Because a subset of large joint-action problems are naturally formulated as multi-agent RL, we additionally evaluated a centralized training, decentralized execution (CTDE) baseline — the predominant paradigm in multi-agent RL (Iqbal and Sha, 2019; Rashid et al., 2020; Yu et al., 2022). As reported in Appendix D, CTDE did not outperform the fully factorized policy and was less stable during training. This suggests that performance is primarily governed by execution-time policy expressivity (i.e., the ability to model coordinated joint-action distributions) rather than by adding global information to the critic.
5.1 State-Independent Sub-Action Dependencies
To evaluate SAINT in environments where sub-action dependencies are primarily state-independent, we use the CityFlow traffic control benchmark (Zhang et al., 2019), where each action corresponds to simultaneous phase decisions across multiple intersections. While coordination is necessary to achieve global traffic efficiency, the structure of these dependencies remains largely unchanged across states.
The results in Figure 2, demonstrate the advantage of SAINT’s set-based action representation. In both environments, SAINT learns faster and achieves higher final performance. The factorized approach is structurally incapable of modeling coordination between intersections and thus performs poorly. The autoregressive model is restricted by its fixed sequential prior, which is misaligned with the unordered nature of traffic-signal control. The flat PPO baseline, forced to learn a monolithic representation without exploiting combinatorial structure, fails to learn an effective policy.
@@ 5.1 State-Independent Sub-Action Dependencies
Figure 2: Learning curves show that SAINT outperforms all baselines in both learning speed and final reward, across both the CityFlow-Linear and CityFlow-Irregular environments.
5.2 State-Dependent Sub-Action Dependencies
We adopt A2C as the standard RL algorithm in this setting. SAINT, the pure factorized baseline, and the AR baseline use identical A2C implementations to ensure a controlled comparison.
Varying Dimensionality

### v3

@@ Abstract
The combinatorial structure of many real-world action spaces leads to exponential growth in the number of possible actions, limiting the effectiveness of conventional reinforcement learning algorithms. Recent approaches for combinatorial action spaces impose factorized or sequential structures over sub-actions, failing to capture complex joint behavior. We introduce the Sub-Action Interaction Network using Transformers (SAINT), a novel policy architecture that represents multi-component actions as unordered sets and models their dependencies via self-attention conditioned on the global state. SAINT is permutation-invariant, sample-efficient, and compatible with standard policy optimization algorithms. In 18 distinct combinatorial environments across three task domains, including environments with $1.35\times 10^{18}$ possible actions, SAINT consistently outperforms strong baselines.  [^1] Code is available at https://github.com/matthewlanders/SAINT
Machine Learning, ICML
@@ 1 Introduction
Traditional RL methods model discrete action spaces with a flat categorical policy, but this becomes intractable in combinatorial settings where the number of actions scales as $\prod_{d=1}^{A}m_{d}$ for $A$ sub-action dimensions with $m_{d}$ choices each. To mitigate this combinatorial explosion, existing approaches (Dulac-Arnold et al., 2015; Pierrot et al., 2021; Tavakoli et al., 2018; Zhang et al., 2018) rely on simplifying assumptions that constrain the representational capacity of the policy class. One family of methods (Tavakoli et al., 2018; Pierrot et al., 2021; Tang et al., 2022; Beeson et al., 2024) factorizes the policy as $\pi(\mathbf{a}\mid s)=\prod_{i}\pi_{i}(a_{i}\mid s)$, which cannot represent interactions between sub-actions. Another class of approaches (Zhang et al., 2018) imposes a fixed autoregressive order, specifying a policy class of distributions of the form $\pi(\mathbf{a}|s)=\prod_{i}\pi_{i}(a_{i}|s,a_{<i})$. This introduces an arbitrary sequence over sub-actions, breaking permutation invariance and impairing learning when the imposed order misaligns with the true dependency structure. Many real-world tasks violate both the independence and fixed-order assumptions. In healthcare, for example, drug combinations can exhibit complex interaction effects — treatments may be safe individually but harmful together, motivating permutation-invariant models for combinatorial action spaces. Our work targets precisely these settings: combinatorial action spaces wherein sub-action indexing is arbitrary or only weakly meaningful, and the fundamental structure lies in sub-action interactions rather than in any prescribed ordering.
@@
More broadly, SAINT reflects a design principle for learning in large combinatorial action spaces: when sub-action indexing is uninformative or only weakly informative, policy classes should be permutation-equivariant while remaining expressive enough to capture dependencies. Commonly used factorized and autoregressive policies encode strong structural assumptions — conditional independence or fixed ordering — that are misaligned with this setting. SAINT instead defines a policy class that is expressive over sub-action interactions, permutation-equivariant by construction, and compatible with standard policy optimization methods.
We evaluate SAINT on challenging benchmark tasks, which exhibit both state-independent and state-dependent sub-action dependencies, including traffic light control (Zhang et al., 2019), navigation (Landers et al., 2024), and discretized DM Control locomotion tasks (Beeson et al., 2024). Our results demonstrate that by learning a more expressive representation of the action space’s internal structure, SAINT consistently outperforms strong factorized and autoregressive baselines, scaling to environments with up to $1.35\times 10^{18}$ discrete actions. Targeted ablations validate the role of state conditioning, demonstrate compatibility with a range of standard online and offline RL learning objectives, and show that the additional cost of modeling dependencies is often offset by substantial gains in sample efficiency. Together, these findings establish that learning explicit representations of sub-action interactions is a practical and scalable approach to decision-making in complex combinatorial domains.
@@ 2 Related Work
Combinatorial action spaces arise naturally in sequential decision problems such as traffic signal control, games, and resource allocation. Prior work has introduced task-specific architectures, imposed domain-specific assumptions, or exploited problem-specific structure (Bello et al., 2016; Chen et al., 2023; Delarue et al., 2020; He et al., 2015, 2016; Nazari et al., 2018; Zahavy et al., 2018; Farquhar et al., 2020). Such methods typically lack generality and require manual design effort. A parallel body of work addresses continuous control problems by discretizing the action space (Barth-Maron et al., 2018; Metz et al., 2017; Tang and Agrawal, 2020; Van de Wiele et al., 2020), which contrasts with our focus on inherently discrete action spaces with combinatorial structure.
@@
A number of general-purpose architectures have been developed to scale RL to large combinatorial action spaces. One strategy reduces complexity by assuming conditional independence across sub-actions (Tavakoli et al., 2018; Pierrot et al., 2021; Tang et al., 2022; Beeson et al., 2024), while another imposes an autoregressive order (Zhang et al., 2018). These approaches improve tractability but either ignore dependencies among sub-actions or introduce arbitrary orderings that break permutation symmetry. Retrieval-based methods such as Wolpertinger (Wol-DDPG) (Dulac-Arnold et al., 2015) scale to large discrete spaces by embedding and pruning candidate actions, but similarly fail to capture the joint structure of unordered sub-actions (Chen et al., 2023). These limitations motivate approaches that can represent dependencies across sub-actions while preserving permutation invariance. In offline RL, SPIN (Landers et al., 2026) follows a representation-first approach: it pretrains and freezes an action-structure model before learning a policy. SAINT, by contrast, is an end-to-end policy parameterization that applies under both offline and online objectives, and can be used either directly as the policy or as a component within modular frameworks like SPIN.
Transformers in Reinforcement Learning
The success of the Transformer architecture (Vaswani et al., 2017) in sequence modeling has led to its widespread adoption in RL. Prominent examples include Decision Transformer (DT) (Chen et al., 2021) and Trajectory Transformer (TT) (Janner et al., 2021), which model trajectories of states, actions, and returns as sequences and autoregressively predict future actions in offline RL. Transformers have also been applied beyond trajectory modeling, including Transformer-based Q-learning in partially observable environments (Esslinger et al., 2022), predictive world modeling (Micheli et al., 2022; Seo et al., 2023), and high-dimensional state representation learning (Kargar and Kyrki, 2021, 2022). In multi-agent and relational RL settings, attention has proven effective for aggregating information across agents or entities, improving reasoning and sample efficiency (Iqbal and Sha, 2019; Zambaldi et al., 2018).
@@
A parallel line of work explores Transformers for action representation. Methods such as RT-1 (Brohan et al., 2022) and RT-2 (Zitkovich et al., 2023) tokenize robot control trajectories and decode actions autoregressively, while Q-Transformer (Chebotar et al., 2023) autoregresses across action dimensions within a timestep. Learned tokenizers like FAST (Pertsch et al., 2025) compress continuous control signals into discrete vocabularies for vision–language–action training, and other extensions interleave state and action tokens or design state-aware tokenizations for multi-component actions (Chen et al., 2021; Shang et al., 2022). While these approaches demonstrate the effectiveness of Transformers for action encoding, they are largely designed for continuous control, offline RL, or VLA settings, and rely on sequential decompositions that impose arbitrary ordering over action components. SAINT instead uses self-attention to model relationships among sub-actions within a single timestep, rather than across time, agents, or external entities. It treats each joint action as an unordered set and applies permutation-equivariant attention (Lee et al., 2019), avoiding the independence assumptions of factored policies and the ordering bias of autoregressive approaches while remaining expressive and tractable.
@@ 3 Preliminaries
Attention is a general computational primitive that allows a model to selectively aggregate information from a set of inputs based on learned relevance scores (Vaswani et al., 2017). Formally, given a set of queries $Q\in\mathbb{R}^{n_{q}\times d}$, keys $K\in\mathbb{R}^{n_{k}\times d}$, and values $V\in\mathbb{R}^{n_{k}\times d}$, the scaled dot-product attention computes output $\mathrm{Attn}(Q,K,V)=\text{softmax}(QK^{\top}/\sqrt{d})V$. Multi-head self-attention extends this by learning multiple independent projections and aggregating their outputs, enabling the model to capture different interactions in parallel. Transformers, which stack layers of multi-head self-attention with feedforward components, have been widely adopted in domains requiring flexible modeling of structured dependencies.
@@ 4.1 State Conditioning
Because sub-action identity alone is insufficient for decision-making, SAINT augments each base embedding $\mathbf{e}_{i}$ with information from the global state $\mathbf{s}\in\mathbb{R}^{d_{s}}$, enabling dependencies to be modeled in a state-aware manner. While several conditioning mechanisms such as cross-attention or concatenation are possible, we adopt Feature-wise Linear Modulation (FiLM) (Perez et al., 2018), which we found to be effective and parameter-efficient (Appendix D). Notably, FiLM preserves the $d$-dimensional width of each sub-action embedding, introducing no additional projection dimensions. Its prior success in incorporating state information in RL (Brohan et al., 2022) further supports this choice.
An MLP $g:\mathbb{R}^{d_{s}}\to\mathbb{R}^{2d}$ processes the global state $\mathbf{s}$ once to produce FiLM parameters $(\boldsymbol{\gamma},\boldsymbol{\beta})=g(\mathbf{s})$, which are then applied uniformly to all sub-action embeddings via an affine transformation:
$\tilde{\mathbf{e}}_{i}=\boldsymbol{\gamma}\odot\mathbf{e}_{i}+\boldsymbol{\beta}\;.$
@@ 4.3 Action Decoding
1:Initialize policy $\pi_{\theta}$ and value network $V_{\phi}$
@@
3:  Collect a batch of transitions $(\mathbf{s}_{t},\mathbf{a}_{t},r_{t},\text{done}_{t})$
4:  Compute $R_{t}$ and $w_{\Phi}(\mathbf{s}_{t},\mathbf{a}_{t},R_{t})$ for each transition
5:  Compute log-probs $\ell_{t}\leftarrow\textsc{LogProbs}(\mathbf{s}_{t},\mathbf{a}_{t})$
6:  Update $\theta$ by ascending $\mathbb{E}_{t}[w_{\Phi}(\mathbf{s}_{t},\mathbf{a}_{t})\cdot\ell_{t}]$
7:  Update $\phi$ by descending $\mathbb{E}_{t}[(V_{\phi}(\mathbf{s}_{t})-R_{t})^{2}]$
@@
11:  Get embeddings $\mathbf{E}\leftarrow[\text{Embed}(1),\dots,\text{Embed}(A)]^{\top}$
@@
14:  Get logits $\text{Logits}_{i}\leftarrow f_{i}(\mathbf{X}_{[:,i]})$ for $i=1,\dots,A$
15:  Compute log-probs for $i=1,\dots,A$:
$\log\mathbf{P}_{[:,i]}\leftarrow\log\pi_{i}(\mathbf{A}_{\text{taken}[:,i]}\mid\text{Logits}_{i})$
@@ 4.4 Compatibility with RL Algorithms
Incompatibility arises when the actor objective requires global operations over the entire combinatorial action space, such as $\mathbb{E}_{\mathbf{a}^{\prime}\sim\pi_{\theta}}[Q_{\Phi}(s,\mathbf{a}^{\prime})]$ or $\max_{\mathbf{a}^{\prime}}Q_{\Phi}(s,\mathbf{a}^{\prime})$. These operations are computationally intractable unless $Q_{\Phi}$ is factorized; however, this changes the structure of the actor target, decomposing it into uncoordinated per-dimension terms and discarding cross-dimensional structure. This breaks alignment with SAINT’s objective of modeling dependencies among sub-actions. SAINT’s learning procedure is provided in Algorithm 1.
@@ 5 Experimental Evaluation
Our experiments evaluate the efficacy of different action representations for modeling complex sub-action interactions across two settings — primarily state-independent sub-action dependencies and state-dependent sub-action dependencies — with results reported in Sections 5.1 and 5.2, respectively. We then evaluate SAINT in the offline RL setting in Section 5.3. Section 5.4 analyzes key architectural choices, quantifying the trade-off between representational power and computational cost, and robustness to Transformer structural parameters.
@@
We compare SAINT to four baselines reflecting the standard representational assumptions for combinatorial action spaces: (1) a factorized policy (Tavakoli et al., 2018; Tang et al., 2022; Beeson et al., 2024), assuming fully independent sub-actions; (2) an autoregressive model (Zhang et al., 2018), imposing a fixed sequential order; (3) Wol-DDPG (Dulac-Arnold et al., 2015), using a continuous embedding; and (4) a flat RL algorithm, which learns a monolithic representation of the full action space without exploiting its combinatorial structure.
@@
These four baselines instantiate the dominant structural assumptions used to scale RL to large combinatorial action spaces. The flat policy corresponds to a monolithic model that ignores compositional structure and treats each joint action as an atomic symbol. The factorized policy enforces independent per-dimension decisions, preventing it from modeling necessary coordination between sub-actions. The autoregressive model imposes a fixed sequential ordering over sub-actions, which can be misaligned with the true, permutation-invariant dependency structure. Wol-DDPG embeds each joint action as a single continuous vector, collapsing the internal structure needed to capture interactions among sub-actions. Our experiments in Sections 5.1-5.3 are designed to test whether these structural assumptions remain sufficient when sub-action indexing is arbitrary or only weakly meaningful, or whether a set-based alternative such as SAINT is required. For completeness, we also evaluate a CTDE baseline in Appendix C; however, it did not outperform the factorized policy, suggesting that performance is limited by execution-time policy expressivity rather than critic information. All results are averaged over five random seeds.
@@
Figure 2: SAINT outperforms all baselines in both learning speed and final reward, across both the CityFlow environments.
5.1 State-Independent Sub-Action Dependencies